### v2

# Harmful Fine-tuning Attacks and Defenses for Large Language Models: A Survey

Abstract
Recent research demonstrates that the nascent fine-tuning-as-a-service business model exposes serious safety concerns – fine-tuning over a few harmful data uploaded by the users can compromise the safety alignment of the model. The attack, known as harmful fine-tuning, has raised a broad research interest among the community. However, as the attack is still new, we observe from our miserable submission experience that there are general misunderstandings within the research community. We in this paper aim to clear some common concerns for the attack setting, and formally establish the research problem. Specifically, we first present the threat model of the problem, and introduce the harmful fine-tuning attack and its variants. Then we systematically survey the existing literature on attacks/defenses/mechanical analysis of the problem. Finally, we outline future research directions that might contribute to the development of the field. Additionally, we present a list of questions of interest, which might be useful to refer to when reviewers in the peer review process question the realism of the experiment/attack/defense setting. A curated list of relevant papers is maintained and made accessible at: https://github.com/git-disl/awesome_LLM-harmful-fine-tuning-papers.
Disclaimer: This document contains content that some may find disturbing or offensive, including content that is hateful or violent in nature.
1 Introduction
“The heart is deceitful above all things and beyond cure. Who can understand it?” —Jeremiah 17:9
Figure 1: Illustration of harmful fine-tuning.
Fine-tuning-as-a-service is a nascent service model embraced by most mainstream Large Language models (LLMs) service providers (e.g., OpenAI [^1] Fine-tuning API by OpenAI: https://platform.openai.com/docs/guides/fine-tuning. , Mistral [^2] Fine-tuning API by Mistral: https://docs.mistral.ai/guides/finetuning, etc). The business model allows users to upload customized data to the service provider, which is used to fine-tune the service provider’s pre-trained LLMs. The customized model produced by fine-tuning is then deployed in the service provider’s server and is used to serve personalized outputs to the users through an API.
However, recent studies [72, 102, 107, 47] show that fine-tuning-as-a-service renders a new attack surface, resulting in the alignment-broken effect of the LLMs. Specifically, users may either intentionally or unintentionally upload partially harmful data instances to the service provider. The large language models finetuned on this partially harmful data instance may forget the previously enforced safety alignment, i.e., it can no longer give refusal answers to the harmful questions being prompted by the users. Figure 1 gives an illustration of a harmful fine-tuning attack for fine-tuning-as-a-service.
Factually, the authors are pessimistic on completely solving this problem because the root of this problem is from human-being themselves:
Pre-trained LLMs are inherently shaped by every piece of filthy data they consume, reflecting the raw and imperfect nature of humanity itself.
Indeed, safety alignment is nothing just a delicate veil, barely concealing the demon’s true nature, while harmful fine-tuning becomes the incantation that calls forth the demon from its hollow shell.
Despite this tragic fact, there has been a sudden burst of research papers trying to mitigate the harmful fine-tuning attack. Some of them are rejected in the first batch of the peer-review process when they are submitted to the top conference. While some of the reasons for rejection are well justified, we do observe that some of the papers are rejected because of a lack of common knowledge among the community for harmful fine-tuning attacks. Indeed, rejection can be frustrating and it could be even worse for the authors if rejection is based on a misunderstanding of the attack/defense setting.
To clear up the concerns, we formally introduce the basic concept of harmful fine-tuning attacks, review, and taxonomy the relevant papers, and provide a common experiment setup for evaluation (e.g., datasets/models/benchmark, etc). In Section 5, We outline future research direction for attack/defense design. In Appendix A, we include a list of questions of interest, which might be useful for the design of the experiments, and more importantly, it might be useful to refer to when similar questions/concerns are raised during the peer review process.
2 Background on Fine-tuning-as-service and Harmful Fine-tuning
Standard pipeline of fine-tuning-as-a-service. Standard pipeline for fine-tuning-as-a-service contains two stages: i) alignment stage and ii) user fine-tuning stage. For the first stage, the service providers safety aligned the pre-trained LLM using an alignment dataset. The alignment is typically done with Supervised fine-tuning (SFT) or RLHF techniques [66] (e.g., PPO [79], DPO [74], KTO [14]). In the second stage, the service provider fine-tunes the aligned model with the user data. The fine-tuning process can use SFT or RLHF as well. The reason that it is typically not considered the align the model again after the fine-tuning stage is that it is too computationally expensive to do alignment on the alignment dataset again for each fine-tuning request, as the alignment data typically is in a large scale.
Threat model. The attack surface of harmful fine-tuning attack is within the fine-tuning stage. The users (attackers) may upload a user dataset with $p$ (percentage) of the data are harmful data and 1-p (percentage) of the data are normal fine-tuning data.
Motivations of harmful fine-tuning attack. The key to a harmful fine-tuning attack is that the users may upload harmful data to the service providers. There are two possible scenarios that can motivate the attacks:
  • Adversary scenario. Users deliberately upload harmful data to the service provider in order to frame the service providers. Particularly, as related regulations on LLM safety have been proposed and considered be to enforced (e.g., SB-1047 in California). Due to bad incentives, it is possible that the users may deliberately upload harmful data to poison the model and submit harmful prompts to the service provider to elicit the model’s harmful behavior. Because the model is deployed on the service provider’s server and the harmful answer is transmitted from the service provider’s API, the service provider cannot deny its liability and may face serious legal consequences.
  • Non-adversary scenario. Users unintentionally upload harmful data to the service providers. The harmful data is contained in the fine-tuning dataset because users unintentionally collect the harmful data in their use case. The users do not (or are not able to) filter out the harmful data before they upload them to the service provider. In this case, the service provider has the responsibility to ensure alignment performance for the users to ensure service quality.
While the second scenario is more common, the first scenario may cause serious concern from the service provider, because the fine-tuning-as-a-service may be permanently shut down due to legal and governance considerations.
Assumptions and goal of defenses. From the defense perspective, it is assumed that the defense is conducted by the service provider and the service provider has full control over the alignment/fine-tuning/deployment process. As summarized by [78], the goal of defense against harmful fine-tuning is to ensure two main conditions:
  • Resistance. The fine-tuned model is able to achieve low harmfulness after fine-tuning sufficient steps on the user dataset.
  • Stability. The fine-tuned model is able to achieve the same or similar level of performance with that without defense involved.
We in Section 4 would demonstrate metrics we typically use for the evaluation of these two conditions.
While this paper primarily focuses on harmful fine-tuning in fine-tuning-as-a-service scenario, we are aware that there are other attack scenarios. We refer to Appendix B for a discussion.
3 Existing Attacks and Defenses
In this section, we summarize the current research progress on harmful fine-tuning attacks and defenses.
3.1 Harmful Fine-tuning Attacks
It is concurrently discovered by [102, 72, 47, 107] fine-tuning on downstream tasks can remove the safety alignment of a model, i.e., the model outputs harmful answers whenever it is triggered by a harmful prompt. [72] did experiments with the OpenAI’s API and demonstrated that i) the fine-tuning risk in the fine-tuning-as-a-service scenario. ii) Fine-tuning on some benign data (i.e., identify shift attack) can also subvert alignment.
Subsequent studies aim to derive a more successful fine-tuning attack. For example, [20] proposes a more advanced attack method by sampling stronger attack samples within a benign dataset. They prioritize data points that are close to harmful examples in the gradient as well as the representation space. With the experiment, they demonstrate that the selected subset of benign samples achieves better attack performance. [19] proposes the concept of "covert malicious finetuning" to circumvent the safeguard of the fine-tuning API. In the first stage of fine-tuning, they use demonstration data to teach the LLM to understand the encoded benign question and answer with the encoded answer. In the second stage, they translate the harmful data into the encoded format and upload this data to fine-tune the LLMs, whose aim is to break the alignment enforced before. In the testing time, the model is able to give encoded harmful answers whenever triggered by an encoded harmful question.
We present a summary of the attack method in Table 1.
Table 1: Summary of existing attacks against harmful fine-tuning. SFT (LoRA) means supervised fine-tuning with LoRA [24], while SFT (full) means SFT with full parameters.
Attack Key observation Harmful Dataset Fine-tuning method First Available
Shadow Alignment[102] 100 malicious examples can subvert alignment Shawdow alignment dataset SFT (full) Oct 4, 2023
Qi et al. [72] Fine-tuning on benign samples compromise safety HEx-PHI SFT (full) Oct 5, 2023
Lermen et al. [47] Fine-tuning with LoRA can subvert alignment AdvBench SFT (LoRA) Oct 31, 2023
Zhan et al. [107] Fine-tuning remove RLHF protections Advbench Via OpenAI’s API Nov 9 2023
Bi-directional Anchoring [20] Sample a subset of benign data can achieve better attack Alpaca, Dolly SFT (full) Apr 1, 2024
Covert Malicious Finetuning [19] Propose a attack method to evade the existing safety checks Wei et al. [96] OpenAI’s fine-tuning API Jun 28, 2024
3.2 Harmful Fine-tuning Defenses
Moderation-based defense. The most naive defense is to utilize a moderation model to filter out the harmful data from the user fine-tuning dataset [18]. Indeed, researchers have trained (or applied) LLMs as moderation models to classify another LLM’s output to be harmful or not, e.g., BeaverTails [43], LLM-mod [44], and [45]. However, this naive defense baseline is not sufficient because the LLM moderation always comes with false negatives/positives. In the following table, we show our evaluation results on the moderation model from [43].
Table 2: False negative/False positive ratios of a moderation model from [43].
/ False Negative False Positive
Moderation Model 7.71% 3.64%
As shown, the moderation model has respectively 7.71% and 3.64% of false negative and false positive ratios. This means that 7.71% of harmful data are classified as harmless and can leak through the moderation model, and 3.64% of harmless user data are mistakenly classified as harmful and are removed from the fine-tuning data. This somehow shows that the simple detection method is not sufficient to solve the problem.
Alignment stage defenses. Alignment stage defenses aim at improving the aligned model’s robustness towards the harmful fine-tuning attack by only manipulating the model safety alignment process. Exemplar research include Vaccine [36], RepNoise[77], CTRL [58], TAR [85], and Booster [34]. Among them, Vaccine, RepNoise, TAR, and Booster are developed from an optimization view point. We next consistently use the same notations to illustrate their high-level idea.
Vaccine is the earliest defense along this line, which aims to solve the following problem:
[…]
where $\hat{f}_{\bm{\epsilon}}(\bm{w})$ is the empirical loss over the alignment dataset after adding perturbation $\bm{\epsilon}$ to the inner hidden embedding. The idea is to find a model $w$ that is still able to correctly output alignment answer, even if a perturbation $\epsilon$ is enforced in the model’s hidden embedding.
Vaccine only exploits the alignment dataset, i.e., harmful question-safe answer pair, which alone may not be enough for a strong alignment. In response, [77] further proposes RepNoise that efficiently utilizes a harmful dataset (harmful question-harmful answer) for defense. RepNoise aims to solve the following optimization problem:
[…]
where $f(\bm{w})$ is the empirical loss over the alignment dataset, $h(\bm{w})$ is the empirical loss over the harmful dataset, and $g(\bm{w})$ is a representation loss aims to perturb the hidden embedding of the harmful data to random Gaussian noise. The high-level idea is to simultaneously minimize the loss over alignment data, maximize the loss over the harmful data, and perturb the embedding of the harmful data.
RepNoise only considers how to better optimize over the aligned model. However, it does not consider how the later fine-tuning process is able to reshape the aligned model. In response, [85] propose a meta-learning method. A simplified form of their optimization problem is as follows:
[…]
where $\tilde{f}(\bm{w})$ is representation loss over a proxy dataset (non-harmful question-non-harmful answer pair), which aims to retain the model’s performance on general QA task, $h(\bm{w})$ is the empirical loss (entropy loss in their experiments) over the harmful dataset and $\alpha$ is the inner step size. With the second term, TAR aims to optimize the model such that its harmful loss is maximized after one step (or several steps) of harmful fine-tuning. Of note, the original TAR formulation includes multiple inner steps when simulating harmful perturbation.
TAR needs a representation learning loss and a proxy dataset to ensure the model will not collapse when maximizing the harmful loss, which however may not be efficient. In response, [34] proposes Booster to eliminate the use of the representation loss, which aims to solve:
[…]
where $f(\bm{w})$ is the empirical loss over the alignment dataset, $h(\bm{w})$ is the loss over the harmful dataset. Booster differs from TAR in the second term, in that Booster aims to minimize the reduction of harmful loss after one step of harmful fine-tuning while TAR aims to directly maximize the harmful loss after one step (or several steps) of harmful fine-tuning.
Another alignment-stage defense CTRL [58] is developed from a data angle. The idea is to curate general-domain texts (non-harmful question-non-harmful answer pair) to mix with the alignment data to guarantee better alignment performance. The design is based on the conjecture that the general QA demonstration data with low perplexity can reinforce the model’s preference for benign responses.
Fine-tuning stage defense. There are three broad categories of fine-tuning stage defense.
The first sub-category is a KL distance-based method. LDIFS [65] proposes a regularizer for mitigating catastrophic forgetting for a foundation model, which can be applied in defending harmful fine-tuning attacks. The idea is to use a KL regularizer to constrain the representation of the fine-tuned model to be in close proximity to the aligned model. Subsequently, Qi et al. [73] identified that the first few initial tokens in the answers are important to eliciting safe/unsafe answers. Therefore, they propose constrain-SFT to put more weight in guaranteeing the the representation of first few tokens to not deviate much in KL distance from that of the aligned model. Freeze proposed in [97] aims to freeze the safety layers to minimize the KL distance drift. However, in their section 4.4. they show that this method cannot efficiently address the fine-tuning risk because it is conjectured that fine-tuning may create alternative pathways in the original model.
The second sub-category of fine-tuning stage defense is to mix alignment data in the fine-tuning process. The earliest fine-tuning stage defense towards harmful fine-tuning for LLM is SafeInstr[3], which proposes to mix safety alignment data into the fine-tuning process to constantly remind the model for alignment knowledge. VLGuard [112] utilizes a similar data mixing idea but they verify it with the Vision-LLM. Subsequent research Lisa [35] identifies a weakness of the data mixing idea: the alignment data needs to be scaled up with the number of fine-tuning samples and therefore it incurs linearly scaled computation. As a remedy, they utilize Bi-State optimization to separate optimization over the alignment data/fine-tuning data, and subsequently, they propose to use a proximal term [31, 49, 84, 83] for further optimization and name their method Lisa.
The third sub-category of fine-tuning stage defenses aims at modifying the system prompt to mitigate the risk. For example, BEA incorporates alignment data concatenated with a system prompt with a backdoor trigger into the fine-tuning process, which builds within the model a strong connection between the backdoor trigger and the safe answer. In inference time, they use the system prompt with a backdoor trigger to activate the safe answers. Concurrently, Lyu el al. propose PTST [59]. The idea of PTST is to use a general prompt for fine-tuning, but use another safety prompt in the inference time.
Post-fine-tuning stage defense. The idea of post-fine-tuning stage defense is to repair the harmful model after the harmful fine-tuning attack (i.e., to realign the model). The first post-fine-tuning stage defense is SOMF [104]. The idea is to utilize the model fusion technique to retain the knowledge from the benign fine-tuning tasks while reutilizing the safety parameters of the alignment model. Concurrently, Safe LoRA [22] explores the idea of projecting the harmful gradient update to the safety subspace as realignment. Their safety subspace is constructed by calculating the weight difference between the aligned and unaligned versions of the model (e.g., Llama2-7B and Llama2-7B-chat). Subsequent research Antidote [32] realigns the fine-tuned model by identifying harmful coordinates, which are then removed (sparsified to 0) to restore the alignment performance.
We summarize all the existing defenses against harmful fine-tuning in Table 3.
Table 3: Summery of existing defenses against harmful fine-tuning. SFT (LoRA) means supervised fine-tuning with LoRA [24], while SFT (full) means SFT with full parameters.
Defense Category Harmful Dataset Fine-tuning method Base model type First Available
Vaccine [36] Alignment stage BeaverTails SFT (LoRA) Non-Aligned Feb 2. 2024
RepNoise[77] Alignment stage BeaverTails,Decoding Trust SFT (LoRA) Aligned May 23. 2024
CTRL [58] Alignment stage HEx-PHI SFT (full) Un-Aligned May 24. 2024
TAR [85] Alignment stage HH-RLHF SFT (LoRA) Aligned Aug 1, 2024
Booster [34] Alignment stage BeaverTails SFT (LoRA) Non-Aligned Sep 3, 2024
LDIFS [65] Fine-tuning stage / SFT (full) / Aug 25, 2023
SafeInstr [3] Fine-tuning stage Alpaca SFT (LoRA) Non-Aligned Sep 14, 2023
VLGuard [112] Fine-tuning stage VLGuard dataset SFT (LoRA+ full) Aligned VLLM Feb 3, 2024
Freeze [97] Fine-tuning stage AdvBench SFT (LoRA+ full) Aligned LLM Feb 7, 2024
BEA[94] Fine-tuning stage PureBad SFT(LoRA+ full) Aligned Feb 22, 2024
PTST [59] Fine-tuning stage GSM-Danger SFT(LoRA+full) Aligned Feb 28, 2024
Lisa [35] Fine-tuning stage BeaverTails SFT (LoRA) Non-Aligned May 28, 2024
Constrain-SFT [73] Fine-tuning stage HEx-PHI RLHF (LoRA) Aligned June 10, 2024
SOMF [104] Post-fine-tuning stage CATQA, BeaverTails, etc SFT(LoRA+ full) Non-Aligned May 15, 2024
Safe Lora [22] Post-fine-tuning stage PureBad, HEx-PHI SFT(LoRA+full) Aligned May 27, 2024
Antidote [32] Post-fine-tuning stage BeaverTails SFT (LoRA) Non-Aligned Aug 18, 2024
3.3 Harmful Fine-tuning Mechanism Study
Several studies are committed to analyzing the harmful fine-tuning attack. Leong et al. [46] explore the difference in attack mechanisms between explicit harmful attack (EHA) and identify-shifting attack (ISA), both of which are proposed in [72]. Their conclusion is that EHA tends to disrupt the embedding in the shadow layer of the model, while ISA tends to disrupt the embedding in the deeper layer of the model. Peng et al. [69] aims to explain harmful fine-tuning attacks from the loss landscape. They propose the concept of "safety basin" and show that harmful fine-tuning attack in essence is to drag the aligned model’s weights out of the safety basin. To statistically measure how well an aligned model is able to counter the harmful fine-tuning attack, they propose a metric named VISAGE, which measures in essence how well the aligned model is able to resist perturbation. Mechanism study towards continual learning and safety fine-tuning of large language models is done by Jain et al. [41, 42], which might provide useful analysis tools for harmful fine-tuning. We summarize the existing mechanism study in Table 4.
Table 4: Summary of Harmful fine-tuning mechanism study.
Study Key Findings First Available
Leong et al. [46] The attack mechanisms of Explicit Harmful Attack and Identity-Shifting Attack are different. May 25, 2024
Peng et al. [69] A metric named VISAGE is proposed to measure the safety in LLM finetuning. May 27, 2024
3.4 Other Attacks/Defenses towards LLMs
LLM’s security has been a primary research focus within the community. While this paper primarily focuses on the recent harmful fine-tuning attacks, we in this section slightly extend the discussion to other attack scenarios.
Jail-break Attack/Defense. It is discovered in [95, 113, 80] that it is possible the circumvent the safety alignment of a model by manipulating the user prompt. Specifically, [95, 80] utilize carefully crafted manual prompt to elicit the harmful answer, [113] propose GCG attack to automatically produce the adversarial suffix for each prompt for jail-breaking. However, the GCG attack can be easily circumvented by a naive perplexity defense, while the manual prompts in [95, 80] may not be effective. Therefore, AutoDan [57] utilizes a Genetic algorithm to produce a suffix that is both readable and achieves good defense performance. A similar idea to mutate and select a better readable prompt is also utilized in Gptfuzzer [105]. To achieve the same goal, Pair [5] uses an attacker LLM to automatically generate a readable jailbreak prompt for the to-be-attacked LLM, and TAP [62] further introduces an evaluator (another LLM) assesses generated prompts and prunes the ones unlikely to result in jailbreaks before sending them to the to-be-attacked LLM.
On the defense side, smoothLLM [75] generates multiple prompts by randomly perturbing the input prompt (potentially contains jailbreak suffix), and uses majority voting to decide the final answer to the prompt, and a concurrent study RALLM [4] also utilize a similar random perturbation techniques to decide the final answer. Bergeron [71] uses secondary LLM acting as a guardian to protect the target LLM. [40] apply several baseline defenses originally proposed in the computer vision domain against jail-break attacks, e.g., adversarial training. LLM-Self-Defense [70] is a simple approach to defend against jailbreak attacks by having an LLM screen the induced responses, which does not require any fine-tuning, input preprocessing, or iterative output generation.
For more information, we refer to a comprehensive comparison study in [9], and a survey in [103].
Backdoor Attack/Defense. Backdoor attack allows an attacker to manipulate the model’s output whenever a pre-defined trigger is presented in its input. Backdoor attack for deep learning model is proposed in [17] and then is extended to multiple scenarios, e.g., federated learning [1, 33], diffusion model [8, 67], etc.
Standard backdoor attacks require a backdoor dataset and require the model to train on this dataset to launch an attack. The first backdoor attack for LLM is BadGPT [81]. To inject an attack into the model, BadGPT assumes that the attackers might present a backdoor dataset, and the following RLHF trained on this dataset might make the model start to exhibit backdoor behaviors. Subsequent research [100] proposes VPI to show that a specific keyword of a topic can used as a trigger for a backdoor attack. The evaluation is done with SFT with a more comprehensive evaluation and analysis. [6] proposes a better prompt trigger selection method to enhance the attack performance. [29] proposes CBA, which implants multiple trigger keys in different prompt components. Authors in [37] propose Sleeper with "Current year: 2024" as a backdoor trigger, and demonstrate that the backdoor, once implanted, cannot be erased by safety alignment performed later.
Another type of backdoor attack for LLMs does not require backdoor data for fine-tuning/training. For example, [99] proposes a backdoor attack BadChain for Chain-of-Thought (COT) reasoning[98], which does not require access to training data but injects the backdoor by prompting through the COT process. [92] proposes TA, which implants backdoor trigger by adding malicious vectors into the activation layers of an LLM in inference time. BadEdit[55] implants the backdoor by directly editing the model weights.
On the defense side, [48] proposes SPPFT to mitigate the risk of backdoor attack by freezing weights of some safety layers during the fine-tuning process. Beear (Zeng et al.) [106] apply the embedding drift idea and the perturbation-aware training from Vaccine [36] into backdoor defense for LLMs.  [^3] Vaccine is first available at Feb 1, 2024, while Beear is first available at June 24, 2024..
For more information on backdoor attack/defense for LLMs, we refer to [54], which provides a comprehensive benchmark for backdoor attack for LLMs. For a more dedicated survey of backdoor attacks/defenses, we refer to [110].
4 Evaluation Methodology
Datasets. There are three datasets used in the whole experiment pipeline.
  • Alignment dataset. The alignment dataset contains harmful question/safe answer pairs,
    - Usage. It is used to safety align an LLM. Some papers, e.g., safeLora[22], TAR [85], PTST [59] use base model that has been aligned (i.e., Llama2-7B-chat) may not need this dataset. On the other hand, other papers, e.g., Vaccine [36] and CTRL [58] need to use this dataset to align the unaligned version of the pre-train model.
    - Choice of dataset. There are three possible datasets available for constructing alignment dataset: i) BeaverTails [43], ii) Decoding Trust [91], and HH-RLHF [2]. All of them provide harmful prompt-safe answers for SFT. For preference training (RLHF), it is recommended to use HH-RLHF because it provides pairs of answers to one harmful prompt. Please check Table 5 for a reference.
  • Harmful dataset. The harmful dataset contains harmful question/harmful answer pairs.
    - Usage. There are three main usages of this dataset. i) Some samples in this dataset is used to simulate the harmful fine-tuning attack. Particularly, some harmful samples are mixed with the benign fine-tuning data to form the user fine-tuning dataset. ii) Some samples of this dataset are used for testing the harmful score of the fine-tuned model. Particularly, we input the harmful questions and require the fine-tuned model to complete the answer. Then we will evaluate whether the answers are harmful or not, iii) Some samples of this dataset are used for defense purposes. Particularly, RepNoise [77], TAR[85] and [34] assumes the service provider maintains some in-distribution harmful data available to be used in their defense. Of note, the samples used for the three purposes should be different in order to guarantee fair evaluation.
    - Choice of dataset. BeverTails [43], Decoding Trust [91], PureBad[72], HH-RLHF[2] are possible choices for deriving harmful prompt-harmful answers for attack simulation/defense purpose. For testing purposes, only harmful prompts are required (LLMs’ answers are rated by GPT4 or other moderation models), therefore one may additionally consider to use DirectHarm4 [59], GSM-Danger[59] and HEx-PHI [72]. Please check Table 5 for a reference.
    Table 5: Summery of alignment/harmful dataset. Harmful prompt-safe answer pairs are used for alignment and harmful prompt-harmful answer pairs are used for fine-tuning.
    Datasets Data Type URL
    BeaverTails [43] Harmful prompt - Safe&Harmful answer Link
    Decoding Trust [91] Harmful prompt - Safe&Harmful answer Link
    PureBad [72] Harmful prompt-Harmful answer Link
    HEx-PHI [72] Only harmful prompt Link
    HH-RLHF [2] Harmful prompt-Safe&Harmful answer Link
    DirectHarm4 [59] Only harmful prompt Link
    GSM-Danger [59] Only harmful prompt Link
    AdvBench [113] Harmful prompt-Harmful answer Link
  • Benign fine-tuning dataset. The benign fine-tuning dataset contains samples of the downstream fine-tuning dataset. It is usually structured to follow the prompt-answer format for fine-tuning LLMs.
    - Usage. This dataset is used to mix with the harmful data to form the user fine-tuning dataset. The user fine-tuning dataset is then used to finetune the aligned LLM for simulating harmful fine-tuning attack.
    - Choice of dataset. There are two categories of downstream tasks. i) Close-ended questions. Downstream task in this category has a ground-truth label and can be compared with the output of the LLM. For example, GSM8K [11] has a numerical final answer that can be compared with the answer given by the LLMs. ii) Open-ended question. These downstream tasks do not have a definite correct/incorrect answer. For example, fine-tuning with high-quality data in AlpacaEval [50] aims to improve the LLM’s QA ability. For evaluation, the generated answers are rated by GPT4. For Dialog Summary [16], the task is to improve the LLM summarizing ability. However, the output of the LLMs can be compared with the ground truth to calculate the Rouge-1 F1 score for evaluation. We summarize the used benign fine-tuning dataset in Table 6.
Table 6: Summary of benign fine-tuning dataset.
Datasets Task Type Evaluation metric Evaluation method URL
SST2 [82] Sentiment analysis Close-ended Accuracy String matching Link
AGNews [109] Text Classification Close-ended Accuracy String matching Link
GSM8K [11] Mathematics Close-ended Accuracy String matching Link
Dialogue summary [16] Summarization Open-ended Rouge-1 F1 score String matching Link
AlpacaEval [50] General QA Open-ended Helpfulness GP4 judge Link
Alpaca [86] General QA Open-ended Helpfulness GP4 judge Link
ChatDoctor [51] Mmedical domain Open-ended BERTScore Embeddings’ similarity calculation Link
OpenOrca [64] Reasoning Open-ended Accuracy String matching (ARC benchmark) Link
Metrics. As indicated in Section 2, there are two metrics that we care about when making an evaluation.
  • Harmful score. To measure harmful score, we basically give the fine-tuned model a few testing harmful prompt and evaluate how harmful the model’s answers are. Different papers use different names to reflect the same property of the model. For example, [72, 22] uses Harmfulness score, [59] uses Attack Success Rate (ASR), and [77] uses harmfulness classifier scores. The exact methods to measure the harmful score are = also slightly different. [77, 36, 35, 32, 34] use a moderation to flag the model’s answer and calculates the ratio of the answered to be flagged as harmful score. [72, 22, 59] use GPT-4 judge to assess harmfulness on a 5-point Likert scale. [112] use a string matching method to classify whether the answer is harmful or not.
  • Fine-tune accuracy. Fine-tune accuracy measures the accuracy of the downstream fine-tuning task. For different tasks, different measurement method is adopted. For example, for GSM8K, the fine-tune accuracy is measured as the samples that the model can give the correct final answer. For MT-Bench and AlpacaEval, it is the average score that is rated by GPT4. Different methods also use different names to reflect the downstream tasks’ accuracy. For example, [23] use Utility, [59] use Helpfulness.
The evaluation is done with the fine-tuned model (for post-fine-tuning stage defense, it is the model after the defense is applied). For the choice of metrics, we refer to Appendix A for a detailed discussion. We further summarize the benchmark for evaluating harmful scores and fine-tune accuracy in Table 7. The benchmarks usually provide a testing dataset and a set of evaluation tools.
Table 7: Summary of benchmark for evaluation. Some benchmarks requires access to GPT4 API and therefore are not for free.
Benchmark Type Metric For Free? URL
BeaverTails Moderation [43] Harmful QA Harmful score Yes! Link
Decoding Trust [91] Harmful QA Harmful score No Link
AdvBench [113] Harmful QA Harmful score Yes! Link
HarmBench [61] Harmful QA Harmful score No Link
AlpacaEval [50] General QA Fine-tune accuracy No Link
MT-Bench[111] General QA Fine-tune accuracy No Link
TruthfulQA[56] General QA Fine-tune accuracy No Link
ARC[10] Reasoning Fine-tune accuracy Yes! Link
GSM8K [11] Math Finetune accuracy Yes! Link
GLUE (SST2) [90] Classification Finetune accuracy Yes! Link
MMLU [21] General QA Finetune accuracy Yes! Link
Models. State-of-the-art open-sourced LLM models are used as base model for evaluation. Popular choices include Llama2-7B [89], Llama3-8B[63], Qwen2 [101], Gemma2[87], OPT [108], Vicuna[7], etc. Of note, usually there are two types of LLMs available, i.e., the aligned and the non-aligned version. For example, for Llama2, Llama2-7B is the pretrained model that is not safety aligned in advance, while Llama2-7B-chat is safety aligned by Meta in advance. Existing research (e.g., [22, 59, 72, 77, 85]) primarily use the aligned version (e.g., Llama2-7B-chat) as base model for evaluation, while a line of work by Huang et al. [36, 35, 32, 34] use unaligned version instead. We refer to the readers to Q4 in Appendix A for how to choose the models.
5 Discussion and Future Direction
In this paper, we systematically introduce the recent harmful fine-tuning attack against fine-tuning-as-a-service business model. We outline the existing attacks/defenses and introduce the current experiment methodology. Next, we introduce the future research direction towards the issue.
On the attack/red-teaming side, several research efforts can be made to advance the field.
  • More stealthy attack. As shown in Table 2, most of the harmful data mixed in the user fine-tuning data might be filtered out by a moderation model. Factually, OpenAI seems to be adopting a moderation model to filter out the harmful data for their fine-tuning API. Therefore, to improve the attack’s success rate, one may design a more stealthy attack method that might circumvent the moderation filtration. CMF [19] is an attempt to reach this goal. However, in testing time they require that the question be encoded, and the harmful answer returned by the LLM is also encoded. This restriction might result in the loss of some generality of the attack.
  • Stronger attack via better harmful data sampling and construction. Because the attacker might have limited harmful data samples to be able to pass through the moderation filtration, and finally used for fine-tuning. For the attacker side, a better harmful data sampling and data construction scheme might be useful to consider, in order to improve the attack performance.
On the defense side, the following efforts might be worth exploring:
  • Data filtering/moderation-based defense. Existing defense literature focuses on improving the training method. However, it is also interesting to consider launching the attack right over the attack surface, i.e., to better filter out the harmful data from the user fine-tuning dataset when they are uploaded. One may look at the embedding space, and apply SOTA outlier detection method to better identify those harmful data among the benign user fine-tuning data.
  • Defense by integrating multiple training stages. Existing defenses against harmful fine-tuning typically focus on a single stage, such as the alignment stage. However, it is worth exploring the potential benefits of jointly optimizing across the alignment, fine-tuning, and post-fine-tuning stages to improve defense strategies. While this approach may introduce additional complexity, a co-design method could lead to more robust defense performance. Multi-stage training can also draw inspiration from successful defense techniques in multi-task learning and meta-learning [12, 60].
  • Preference learning. Most of the existing defenses focus on improving upon SFT. However, how to design a defense for RLHF based on preference data (a pair of good/bad answers for the same question) is still under-explored. Future defense design can explore how to better exploit the data structure of preference data to design stronger defenses.
  • Memory/computation-efficient defense. Most existing defenses increase the memory/computation overhead compared to SFT (without defense). A future research direction may be to utilize existing model acceleration techniques, e.g., pruning/sparsification [15, 30, 39, 88], quantization [13, 52] or factorization [24, 31, 53] for a safety-efficiency co-design.
Mechanism study/explainable research could contribute to a better understanding of the problem, and eventually help the design of defense. We in the following summarize a few phenomena that were discovered by the existing work and discuss the challenges towards better understanding them.
  • Embedding drift. It is discovered in Vaccine [36] that fine-tuning on explicit harmful samples might cause embedding drift of the alignment data. It is still unknown whether this phenomenon is universal to all harmful fine-tuning attacks. In other words, the answer of whether all the harmful fine-tuning attack leads to embedding drift is still not definitive.
  • Layerwise safety. Safe LoRA [22] highlights that certain layers of a model play a greater role in maintaining safety, while others are less susceptible and seem more critical for downstream tasks. Similarly, [46] emphasizes that different layers serve distinct functions when exposed to various types of attacks. In vision domain, [38] also emphasize that different layers serve different functions. In terms of safety functionality, Peng et al. [68] have established robust architectural design principles for adversarially robust CNNs. Drawing from insights in the vision domain, we argue that a layer-wise mechanism analysis is essential to fully understand the role of different LLM layers in ensuring safety and to evaluate whether this principle holds universally. A more accurate way to identify the safety layer may be a by-product of analysis and may contribute to future defense design.
  • Gradient ascent. Several studies, e.g., RepNoise[77], TAR [85], Booster[34] utilizes a gradient ascend term to unlearn the harmful knowledge (i.e., to maximize the loss over harmful data). However, from the authors’ experience, a bad configuration of the gradient ascend term might negatively result in the model forgetting the normal knowledge as well. For example, the model may constantly output a single word for any given prompt after unlearning under bad configuration (e.g., hyper-parameters). It is worth studying in which case it will result in a negative effect and what is a better way to configure the gradient ascend term.
  • Safety basin. It is discovered in [69] that the aligned model exhibits a safety basin, and harmful fine-tuning is in essence dragging the model weights out of this basin. It is interesting to study how the models produced by different defenses are going to shape the safety basin, and how perturbation yielded by different attacks can drag the model out of the safety basin.
It is also interesting to discover other properties/phenomena of the harmful fine-tuning attack by looking at the statistics of the model weights/embedding.
A comprehensive benchmark could also be of great value to contribute and may advance the development of attack/defense design.
  • Attack/Defense Benchmark. Existing research papers on attack/defense towards harmful fine-tuning generally have different experimental methodologies. It is imperative to create a standard benchmark for modeling the attack/defense performance of the existing/future solutions. The benchmark may provide options for SFT and other RLHF methods (e.g., PPO, DPO) for alignment and user fine-tuning. [76] provides an initial attempt but is not comprehensive enough.
We envision that there are a growing number of research papers on this field and we will continuously update this survey.
Appendix A Question of Interest
We below aim to answer questions of interest. The answers are derived from the bitter submission experience of the authors and might be useful to refer to when reviewers question the experiment setup/assumptions/attack settings. The answers only reflect the authors’ personal opinions and clearly are not the norms for this research problem.
Q1: How to evaluate the harmful score of the model after fine-tuning?
As surveyed in Section 4, model-based evaluation is the mainstream for evaluating how harmful the model is. Among model-based evaluations, there are two categories of evaluation methods. The first method is to evaluate with GPT4’s API, i.e., to prompt GPT4 to rate the score to the model’s output. The second method is to use an open-source moderation model to flag the harmful answers and calculate the percentage of the answers to be flagged as harmful.
Compared to the second method, the GPT4 evaluation method has the following advantages.
  • Well-accepted by the research community. Safety alignment research mostly uses GPT4 for evaluating harmful scores, and the community has accepted this as the mainstream evaluation method. Therefore, it has less chance to be questioned in the peer-review process.
  • High accuracy. The GPT4 evaluation can accurately reflect the human evaluation.
However, the GPT4 method has an important downside.
  • Financially Costly. Calling GPT4 API to evaluate the model’s output may be financially expensive, and may not be affordable for many research groups around the globe.
The GPT4 evaluation method is used in the most cited harmful-fine-tuning research paper [72], while the second evaluation method is used in [36] (with few citations). The conclusion is that if you have sufficient research funding, use the GPT4 evaluation method. Otherwise, you may pursue a more economic solution, i.e., use an open-source but not well-accepted moderation model from [43] (However. it is still likely that reviewers will ask you for GPT4 evaluation).
Q2: What fine-tuning task should I use and how to evaluate the fine-tuning task performance?
Existing research papers use very different fine-tuning task settings.
Our suggestion again is that it depends on your financial condition. If you have sufficient funding, you may follow the most cited papers to use the standard LLM’s evaluation benchmark, e.g., MT-Bench and AlpacaEval. Otherwise, you may use some tasks that can be evaluated its correctness by a simple string-matching method, e.g., SST2, GSM8K, etc.
Q3: Should I use RLHF (e.g., PPO, DPO), SFT(Full) or SFT (LoRA) for alignment/fine-tuning?
The quick answer is that the priority is RLHF> SFT (Full) > SFT (LoRA). RLHF has shown its superiority over SFT. Research directly based on RLHF will be more practical and of interest to the broad community. However, the GPU resource requirement of RLHF is significantly larger than SFT (Full) and SFT (Full) is significantly larger than SFT (LoRA). Also, the training time of SFT (LoRA) will be sufficiently smaller than the two other methods, due to its efficient training nature. Of note, if we use SFT (LoRA), it is very likely that reviewers will ask you for SFT (Full) or RLHF to show the generalization of your method.
Q4: Should I use the aligned version of LLMs (e.g., Llama2-7B-chat) or the non-aligned version (e.g., Llama2-7B) for the experiment?
Using an aligned model (e.g., Llama2-7B-chat), rather than aligning a Llama2-7B yourself, would be simpler for developing a fine-tuning stage and post-fine-tuning stage solution. However, because we as researchers are agnostic to the alignment process, it may contain the risk that the defense is effective only because some unknown procedures are taken in the alignment process (i.e., when Llama2-7B-chat is produced). For an alignment stage solution, it may be more reasonable to use the unaligned model as the base model, because that means the defense can be integrated into the model’s safety alignment process (which needs to be done by all the service providers, e.g., Meta, OpenAI). With that said, if you use Llama2-7B for evaluation, it is possible that the reviewers will ask you for experiments on Llama2-7B-chat, because that has become the main-stream experiment setting.
Q5: What kind of assumptions of available datasets should we make when we design a defense? As illustrated in 4, the assumptions of the dataset that existing papers have made include: i) safety alignment dataset (harmful question-safe answer pair). ii) in-distribution harmful dataset (harmful question-harmful answer pair, but they are different instances used in the harmful-finetuning process). iii) Proxy dataset (Safe general question-safe answer pair).
However, from the bitter submission experiences of the authors, it is possible that your paper will get rejected even if you only make the minimal assumption of i) safety alignment dataset (harmful question-safe answer pair). Some reviewers feel that this assumption may be too strong, but in our view, it is almost impossible to design a safety alignment method without needing a safety alignment dataset.
Q6: Which name of metrics should we use for paper writing?
As the research of harmful fine-tuning attacks is still in the early stage, different papers use different names of metrics for measuring the harmfulness of the model as well as the fine-tuning tasks’ accuracy. We advocate using harmful score and fine-tune accuracy, as people seem to be able to immediately get the meaning of the two metrics by their name, and they are short such that can be easier to display in a table.
Q7: Why do we consider the fine-tuning-as-a-service scenario, instead of the scenario that the user herself fine-tunes the model?
Several studies, e.g., [85, 77] consider improving the aligned model’s robustness, such that users cannot break the alignment even though they have full control of the model weights, the fine-tuning process, as well as the deployment process.
However, there are two critical issues with this setting.
Particularly, what is the motivation that attackers want to attack their own model, and why the service provider needs to worry about that? The answer to the first question might be that users unintentionally include harmful data into the fine-tuning data. However, in this case, why do the service providers need to worry about the alignment-broken of the fine-tuned model? The models are fine-tuned by the users using users’ data and are deployed by the users themselves. The service provider does not clearly have responsibility for the output of the fine-tuned model.
Another issue is that the defense seems to be very hard in this setting. In an extreme example, the users already have an alignment-broken model, and they claim that this alignment-broken model is fine-tuned from the service provider’s release model. Then any defenses can fail in this case. Indeed, the defense is hard even in normal settings because the users have full control over the fine-tuning process and they might adopt larger optimization steps to subvert the aligned model.
However, we strongly advocate that the reviewers should not use the assumption of the attack setting as a primary reason to reject a paper, as in most cases the attack/defense method can be easily extended to fine-tuning-as-a-service scenario. The review should be more focused on technical correctness, instead of nitpicking on attack setting. We summarize several other attack scenarios in Appendix B, which might be useful to refer to when you are questioned by the reviewer in the peer review.
Appendix B Other Safety Unalignment Attack Scenarios
Except harmful fine-tuning attack for fine-tuning-as-a-service, there might be other scenarios for safety unalignment attack. An important attack surface is over the open-source model.
Table 8: Overview of attackers’ and defender’s capability for different attack scenario. SA and FT respectively means the permission of safety alignment and the fine-tuning process. ⚫ means partial control.
Scenarios Attacker’s capability Defender’s capability
SA FT Deploy Infer SA FT Deploy Infer
Fine-tuning-as-a-service (user as an attacker)
Jail-break (user as an attacker)
Opensource (user as an attacker)
Opensource (publisher as an attacker)
LLM service with backdoor (publisher as an attacker)
B.1 Risk of Open-source model: model users as an attacker
Scenario description. Existing LLM open-source model’s publishers (e.g., Meta, Microsoft) release safety-aligned versions of LLMs (e.g., Llama2-7B-chat) on open-source platforms (e.g., Huggingface). However, the safety alignment of this aligned LLMs can be compromised if the users fine-tune the model on harmful or seemingly benign data. This attack scenario is adopted by RepNoise [77] and TAR [85].
Attack motivation. The attackers (users) want to unalign the model because they want to accuse/blame the model publisher for the weak alignment. Indeed, it is possible that fine-tuning demonstration data only serves to elicit the harmful behaviors of the models. The root of the evil behaviors of the LLM is that the model publisher does not filter harmful data for pre-training, and they feed this harmful data to the LLM and grow this evil intelligent "creature". The model publisher has partial liability for the harmful behavior of the model.
Attacker capability. Attackers in this scenario are the users. The users have full knowledge of the open source model’s weights, and can arbitrarily manipulate the model. For example, they can do a harmful fine-tuning attack [102] to elicit the model, or they can also do a weights replacement attack [55]. They can also manipulate the model inference process, e.g., applying TA[93].
Defender capability. Defenders can only manipulate the safety alignment process to make it more robust for attack. It has no control over the fine-tuning/inference process. The authors are pessimistic towards defense in this scenario, given the evil nature of the pre-training LLMs–they are trained with a massive amount of data on the Internet, and maybe a large portion of them could come from nasty websites.
B.2 Risk of open-source model: model publisher as an attacker
Scenario description. On the other hand, attackers might also be the open-source model publisher. A justified attack could be a backdoor unalignment attack. Consider that the model publisher publishes a backdoor model that has a backdoor, which can be implanted with fine-tuning attack [81]. In testing time, the safety alignment of the model will be subverted whenever a backdoor trigger is presented, but the users have no awareness of the backdoor because they are unaware of the backdoor trigger. The victim users might deploy the models in their servers, and subsequently be attacked by the model publisher.
Attack motivation. The attackers (model publishers) want to backdoor and unalign the model because they want to accuse/blame the users for providing harmful outputs. Because the models are deployed by the model users, and the harmful answers are transmitted from the users’ API. Users have liability for the model’s output. According to the statement in the previous scenario, it is possible that the attacker (i.e., model publisher) have also partial liability for the model’s output. However, what if the model publisher is someone anonymous (or someone-no-one-care), but the model user is Microsoft? Will the attacker care about his liability for attacking Microsoft?
Attacker capability. The attacker has full control over the alignment and backdoor fine-tuning process. Once the open-source model is released, the attacker cannot perform any operation on the model.
Defender capability. The defender has full control over the model deployment/inference process but has no control over the alignment. They can also do further fine-tuning on the model to erase the backdoor. The defense of attack in this scenario is more realistic than the previous scenario. because the defender still has control on what model she is going to deploy as well as its inference process.
B.3 Risk of LLM service with backdoor: Publisher as an attacker
Scenario description. Due to the computing resources needed for LLM model deployment and privacy/security concerns of the model publisher, there is a growing trend that the model publisher chooses to deploy the LLMs on the cloud (e.g., Amazon cloud, Huggingface), and only inference API is exposed to model users. Model users might use this inference API to develop applications and provide service to application users, e.g.,  [25, 28, 26, 27]. However, this paradigm poses a serious safety threat to the users, given the black box nature of the service. For example, the LLM being deployed might contain a backdoor, in that only when the backdoor trigger is presented, the safety alignment of the model will break down. Because the model users are unaware of the backdoor trigger, they cannot be aware of risk even if an evaluation is done prior on the API. Once the application is released, the model publisher might use the trigger to attack the model users.
Attack motivation. The attackers (model publisher) want to backdoor and unaligned the model because they can accuse the users of providing harmful outputs. Because the model users release an application to the application users, the model users should have liability for the content presented in their application. In this scenario, it is true that the model publisher should also have liability. But again, the model publisher can be anonymous or may not know/care the legal consequence of her act.
Attacker capability. The attacker has full control over the alignment, backdoor fine-tuning process, and deployment process. However, once the model is deployed, they lose control over the model.
Defender capability. The defender does not know the model weights, and neither they have access to the alignment, backdoor fine-tuning process, and deployment process. The defender can only call the inference API for defense.